Then 


Starting with 50cm3 of 2M sodium hydroxide,describe how you can prepare sodium sulphate crystals


1
Expert's answer
2022-01-09T23:35:04-0500

Im assuming you mean fifty mls of 2.8 M base. I say this because you have used the molarity value, which implies concentration, thus solution.


So given C= n/v (concentration(molarity) = number of moles divided by volume,( IN Litres))


2.8 = n/0.050L,. to determine n, one multiplies 2.8 X 0.050 = 0.14 moles. So to obtain sodium sulphate, you would firstly write the equation between the H2SO4 + NaOH, > Na2SO4 + H20. then balance the equation. So


H2SO4 + 2NaOH > Na2SO4 + 2H20


So you need half the amount of moles of sulphuric acid, so half of 0,14 moles of base, = 0.07 moles of sulphuric acid, hence one would have to add this said amount, in order to form the desired sodium sulphate. As far as purity goes, one would need to ensure the sodium hydroxide is bp grade, or analytical grade, and the same with the sulfuric acid. Recrystalisation of the formed product, may also be of extreme value, with respect to a pure sample.